OpenStudy (anonymous):

Which expression is the simplified form of the square root of k to the twenty-first power.? A:k5 B:k10the square root of k. C:k10 D:k4the square root of k.

OpenStudy (helder_edwin):

\[ \LARGE \sqrt{k}^{21}=\sqrt{k}^{20+1}=\sqrt{k}^{20}\sqrt{k}=? \]